OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 −1 0 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (austinl):

Ok, what I would do is solve one of the equations for one variable and then plug it into the other equation. Do you think you could do that?

OpenStudy (austinl):

Actually, correction, I would solve for y in one and plug it into the other.

OpenStudy (zzr0ck3r):

I would say substitution but that is going to be fractions so try this 2x − 3y = 3 5x − 4y = 4 multiply the first equation by -4 and the second by 3 -8x+12y=-12 15x-12y=12 add the two equations to cancel the y terms 7x=0 so what is x?

OpenStudy (anonymous):

0

OpenStudy (zzr0ck3r):

correct ,do you see how I did it?

OpenStudy (anonymous):

yes, thank you! can you help me with one more?

OpenStudy (zzr0ck3r):

I made it so that the y terms would cancel when we add the equations...

OpenStudy (zzr0ck3r):

close this and post another question.

OpenStudy (austinl):

The way I did it is the way that you avoided, but I personally love fractions as they are fun to me. All the same, I arrived at the same conclusion.

OpenStudy (zzr0ck3r):

yeah I like em as well, but people like us are few and far between:)